a 0.02M solution of latic acid has ph =2.8 calculate the approximate ka for lactic acid

Grade:12th pass

1 Answers

Sahani Kumar
5 years ago
Given: c= 0.02   pH=2.8
 Now, pH= -log[H+]            [H+]= concentration of H+
           2.8= - log[H+]
           [H+]=1.58*10^-3
We know that α =[H+] = 1.58*10^-3
     Now, Ka=c α^2
               Ka= 0.02 * (1.58*10^-3)^2
               Ka=5*10^-8
Regards 
Askiitian member
If you like this answer please encourage us by approving this answer.